OpenStudy (wintersuntime):

Can someone be nice enough to help me with this question .>> 3x-10=2 what's x ?

OpenStudy (anonymous):

x = 4 3 x 4 = 12 - 10 = 2

OpenStudy (anonymous):

3x=12 x=12/3 x=4

OpenStudy (wintersuntime):

yes I was right thanks

OpenStudy (wintersuntime):

need help on one more

OpenStudy (anonymous):

Yup :) Ask away

OpenStudy (wintersuntime):

3x+6=x+2

OpenStudy (anonymous):

3x=x-4 2x=-4 x=-4/2 x=-2

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

Simplifying 3x + 6 = x + 2 Reorder the terms: 6 + 3x = x + 2 Reorder the terms: 6 + 3x = 2 + x Solving 6 + 3x = 2 + x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1x' to each side of the equation. 6 + 3x + -1x = 2 + x + -1x Combine like terms: 3x + -1x = 2x 6 + 2x = 2 + x + -1x Combine like terms: x + -1x = 0 6 + 2x = 2 + 0 6 + 2x = 2 Add '-6' to each side of the equation. 6 + -6 + 2x = 2 + -6 Combine like terms: 6 + -6 = 0 0 + 2x = 2 + -6 2x = 2 + -6 Combine like terms: 2 + -6 = -4 2x = -4 Divide each side by '2'. x = -2 Simplifying x = -2